# Huge Ensembles Part II: Properties of a Huge Ensemble of Hindcasts Generated with Spherical Fourier Neural Operators

### v2

@@ 3.1 Sampling the Forecast Distribution with Huge Ensembles
For all statistics, smaller ensemble sizes result in large uncertainties and, in some cases, large biases relative to the full ensemble. Estimates of the HENS mean are unbiased (i.e., the empirical bias is near zero) with as few as 10 members and the HENS standard deviation with as few as 200 members. However, these estimates are associated with large sampling uncertainty, well in excess of six times the uncertainty of the full ensemble. Larger ensembles are necessary for unbiased estimation of the 10 and 90 percentiles, on the order of . For ensembles smaller than 1000, the sign of the bias is notable: estimates of the 10 and 90 percentile are too large and too small, respectively, highlighting the under-sampling of even these moderate percentiles for smaller ensembles. Even with 1000 members, there is still sampling uncertainty associated with moderate percentiles, nearly three times that of the full ensemble, and larger ensembles are necessary to reduce this uncertainty. For the most extreme percentiles, representing 1000-day events, nearly the full ensemble is needed to obtain empirical estimates that are unbiased. In Section F.4.2, we present extreme value methods to calculate the 0.1 and 99.9 percentiles based on an ensemble with a given number of members. Compared to directly calculating the percentiles from the given ensemble, extreme value theory leads to better estimation of the extremes (Figure 4): estimates are unbiased for ensembles as small as 3000, which is a significant improvement relative to empirical estimates which require . However, as with the 10 and 90 percentiles, for smaller ensembles both the empirical and extreme value theory estimates are not extreme enough, again illustrating that smaller ensembles do not properly sample the extreme tails of the distribution.
@@
Across all statistics considered here, larger ensembles lead to significantly more confident estimation of each property of the large ensemble. This is a key value-add of HENS: it enables confident characterization of both the mean and extreme statistics of the forecast distribution, and it quantifies the uncertainty associated with smaller ensemble sizes. Using this information, the users can select a desired ensemble size based on their use case and an acceptable level of uncertainty.
@@
While Figures 2 and 4 characterize the global land mean of each ensemble member, we present a demonstration of the ensemble forecasts at a specific location during a heatwave in the USA Midwest. Kansas City, Missouri, USA had a significant heat-humidity event on August 23, 2023 at 18:00 UTC. According to ERA5, the 2m temperature and dewpoint reached 307 K and 298 K, respectively. A key benefit of our SFNO configuration is that it includes 2m dewpoint temperature as a prognostic variable. To consider the combined effect of both temperature and humidity at the surface, we calculate the heat index introduced by Lu and Romps (2022). This heat index has been updated to account for particularly hot and humid events, during which the standard heat index is undefined. At this time, the heat index in Kansas City was 316 K (43 C).
@@
Figure 5a shows the HENS forecasts of the heat index, as a function of initial time. Despite the significantly larger size of HENS, its ensemble range becomes narrower with lead time. Therefore, HENS does not overpredict extreme values for its forecasts at all lead times; its forecasts still have coherent spread as a function of lead time. At a 10-day lead time, both IFS and HENS predicted a warmer than average temperature. However, the verification air temperature and dewpoint temperature lies at the tail of both ensemble’s forecast distributions (Figure 5b). At a ten-day lead time, none of the IFS members successfully capture the magnitude of both 2m air temperature and 2m dewpoint temperature. However, HENS does include members that capture the simultaneous intensity of both these values (Figure 5b). With the large sample size from HENS, it would be possible to study the precursors, drivers, and statistics of the observed extreme. The huge ensemble can more accurately sample the tails of the distribution, so at a 10-day lead time, it includes ensemble members that are sufficiently hot and humid to capture the observed extreme. Figure 5 uses the same method to visualize large ensembles Li et al. (2024), though it uses 2m dewpoint temperature instead of total column water vapor.
@@
The HENS ensemble also enables exploration of counterfactual realities. If the forecast is reliable, it can be used to quantify the probability of the extreme event as a function of all possible outcomes. Some HENS members projected that the heatwave could have been warmer and drier, with a higher air temperature and a lower dewpoint temperature (Figure 5b). Alternately, other HENS members indicate that the heatwave may have not occurred altogether, as they predict 2m temperatures near or below the climatological mean. These different counterfactuals can result in different climate impacts: hot and humid extremes can be particularly challenging for human health, while hot, dry extremes can have adverse impacts on crop yields and create conditions conducive to wildfire spread. IFS does not contain any sampling of the counterfactual hot, dry version of the event beyond 311 K air temperature and 291 K dewpoint. Because HENS thoroughly represents the forecast distribution of future outcomes, one could characterize the dynamical drivers and physical processes associated with each outcome. In particular, HENS could enable a clustering analysis of the drivers and large-scale meteorological patterns that would result in the observed hot, humid heatwave; a hot, dry heatwave; or no heatwave at all. Furthermore, the statistical likelihood of each cluster can be characterized in light of all possible future outcomes. Thus, HENS provides a thorough representation of internal variability, and it is a valuable tool for studying the climatic statistics of extremes.
@@ 3.2 Sampling the Observed Distribution with Huge Ensembles
At each grid cell and time in summer 2023, we convert the ERA5 value into its Z score. Figure 6 shows that HENS provides large sample sizes for the ERA5 events that occurred in summer 2023. In the majority of cases, HENS includes multiple ensemble members that simulate an event that is at least as extreme as the verification value in ERA5. For even the most rare events that are 4 standard deviations away from their climatological mean, HENS usually includes at least samples of events with at least that magnitude. For events that are 2 and 3 standard deviations away, there are hundreds of ensemble members in HENS that meet or exceed the ERA5 value. We note that there are very few instances where HENS misses an event entirely: we quantify this occurrence in Section 4.
@@
The second requirement for calculating robust statistics is that some ensemble members should accurately simulate the true event. This is necessary to calculate the probability of the observed outcome compared to other possible outcomes. To assess the suitability of HENS for this requirement, we calculate the RMSE of the best ensemble member. Buizza and Palmer (1998) originally introduced this metric for ensembles with up to 32 members. At each grid cell, they choose the ensemble member with the smallest RMSE across all members. Across multiple forecasts and lead times, they quantify the RMSE of this best member.
@@
This metric is actively used in the study of extreme weather events. Using the most accurate ensemble members, recent studies identify possible drivers for extreme weather events. In an operational ensemble weather forecast, Mo et al. (2022) identify the members that had the most accurate forecasts of the 2021 Pacific Northwest Heatwave. These best members correctly forecast the extent and inland location of a warm-season atmospheric river, which served as a source of latent heat for the heatwave. Leach et al. (2024) also use this metric. They study the ensemble member that is nearest to the observed temperatures. They show that ensemble members that predicted warmer temperatures were associated with low cloud cover and a high geopotential height anomaly. With HENS, these types of analyses can be conducted at scale. Because of its large ensemble size, HENS reduces the RMSE of the best ensemble member. By considering these ensemble members with the highest skill, it is possible to analyze the drivers of specific events. Compared to smaller ensembles, HENS more thoroughly covers the space of future outcomes.
As a function of ensemble size, this RMSE of the best ensemble member systematically decreases (Figure 7). With larger ensembles, there are more members that closely approximate the observed value. This opens the door for future analysis to determine why these members correctly forecast the event and to validate whether they forecast the right value for the right reasons. Additionally, the statistics of the best ensemble members can be compared with the full spread of the ensemble to assess the likelihood of the event. We note that the minimum ensemble RMSE metric is not useful in making weather forecasts: while huge ensembles virtually guarantee that at least one member will reasonably match the observations, there is no way of knowing ahead of time which member that will be.
@@ 4 Validating Huge Ensemble Weather Forecasts
4.1 Sampling the bulk distribution and the conditional distribution
Next, we discuss opportunities for using HENS to improve extreme weather forecasts. The CRPS is an overall summary statistic for evaluating ensemble forecasts. Allen et al. (2023); Gneiting and Ranjan (2011) discuss variations of CRPS that focus on extreme weather and the tails of the forecast distribution. One variation is the threshold-weighted CRPS (twCRPS). twCRPS is calculated analogously to CRPS except all values of the ensemble forecast below a pre-specified threshold are converted to the threshold itself (Allen et al., 2023). In Part I, we showed that SFNO-BVMC and IFS have comparable CRPS and twCRPS scores. In this section, we discuss the effect of ensemble size on twCRPS and CRPS, and we go beyond these metrics to directly consider the tail of the forecast distribution.
Because of HENS’s ability to sample low likelihood events in the forecast distribution, we assess the CRPS values of HENS and the 58 member SFNO-BVMC from Part I in Figure 8. For this section, we use the 58-member ensemble as a benchmark ensemble, because it is the standard configuration used in Part I for all diagnostics. It uses all 29 checkpoints, with 2 initial condition perturbations per checkpoint (one bred vector added to and subtracted from the initial condition). Based on the CRPS and twCRPS scores, HENS is slightly better than the 58-member ensemble due to its larger ensemble size, but these scores do not show large percent changes for HENS compared to the 58-member ensemble (Figure 8a and b). The CRPS and twCRPS scores improve by less than five percent. These scores have mostly converged for ensembles with 58 members.
@@ 4.1 Sampling the bulk distribution and the conditional distribution
CRPS and twCRPS compare the CDF of the ensemble forecast to the verification value. See Equations 1 and 3 of Part I for the equations for CRPS and twCRPS. CRPS and twCRPS scores are calculated based on CDFs that use all members of the ensemble forecast. CRPS uses all members of the raw forecast distribution, and twCRPS uses all ensemble members, but the ensemble is transformed such that all members below the threshold are set to have a value at the threshold. The Dvoretzky–Kiefer–Wolfowitz (DKW) inequality (Massart, 1990) quantifies the maximum difference between a true population PDF and an empirical CDF constructed from samples. Here, we use the HENS ensemble forecast as the population CDF, and the 58-member ensemble would be the sample empirical CDF. The difference between these two CDFs scales with a factor of (Section G). Therefore, we expect the 58-member ensemble CDF to closely approximate the HENS CDF, which uses 7,424 members. Based on this argument, we also would not expect the Extreme Forecast Index to change considerably when using HENS, instead of the 58-member ensemble. Equation 2 of Part I shows that the EFI calculation also heavily relies on the ensemble CDF calculation.
@@
A key value-add of HENS is that it provides more ensemble members to calculate the ensemble’s conditional distribution, given that the forecast is above the threshold. This conditional distribution is created by truncating the ensemble forecast distribution at the 95th percentile. When calculating the CRPS using this conditional distribution, the score is called the outcome-weighted CRPS (owCRPS). Mathematically, owCRPS is
@@
where is a random variable from the ensemble CDF. We use the weighing function , based on a threshold . Here, we use the 95th percentile of 24-year ERA5 climatology as the threshold.
@@
We show the CDF of HENS and the 58-member SFNO-BVMC in Figure 9a. Due to its smaller size, there is more noise in the 58-member CDF compared to the HENS CDF. In Figures 9b, c, and d, we compare the distributions that are used to calculate CRPS, twCRPS, and owCRPS. These three scores use the raw ensemble distribution, the transformed ensemble distribution (described above), and the conditional ensemble distribution, respectively. twCRPS results in a point mass being placed at the threshold (Figure 9c), since all ensemble members below the threshold are converted to having a value at the threshold itself. owCRPS is created from the distribution truncated at the threshold.
CRPS and twCRPS show small percent improvements with larger ensemble size. This is because they use the entire ensemble to calculate the score. On the other hand, owCRPS uses the conditional distribution; it constructs the CDF only from the members above the threshold. Figure 1 of Allen et al. (2023) provides a demonstration of calculating CRPS, twCRPS, and owCRPS using an idealized Gaussian. We mirror their demo in Figures 9b, c, and d using real HENS forecasts at a sample location in Shreveport, Louisiana.
@@
In Figure 9d, the 58-member ensemble had 10 members above the threshold, and HENS had 1340 members above the threshold. While twCPRS and CRPS use all the ensemble members, the owCRPS score only uses the members above the threshold. In this instance, 10 members do not adequately characterize the tail of the forecast distribution (Figure 9d). The small ensemble has a significantly different conditional distribution than HENS’s conditional distribution. For the smaller ensemble, the owCRPS score has not yet converged as a function of ensemble size, and HENS shows a marked improvement in owCRPS: it reduces the CRPS from 0.63 K to 0.5 K (a relative reduction of about 20%). Unlike CRPS and twCRPS, the number of ensemble members used in the owCRPS varies for each forecast at each grid cell: it depends on how many members are above the threshold.
@@
The owCRPS is calculated only when the extreme event actually occurs, so it is not a statistically proper scoring rule. Proper scoring rules are minimized when the forecast distribution matches the distribution from which the observation is drawn (Gneiting and Raftery, 2007). They cannot be hedged by overpredicting extremes (See Part 1 for a discussion on the relationship between CRPS and proper scoring rules for extreme forecasts.) If a system overpredicted extremes, it would falsely appear reliable. Since the owCRPS is only calculated when extremes actually occur, the forecast would falsely appear reliable (Lerch et al., 2017). This is the essence of the forecaster’s dilemma (see Part I for a deeper discussion). However, HENS does not appear to be overpredicting extremes, which could result in an better owCRPS score. HENS has a comparable (and slightly better) CRPS and twCRPS as the 58-member ensemble. If HENS got an improved owCRPS by overpredicting extremes, then the CRPS, twCRPS, reliability, and the spread-error ratio would significantly degrade. In Figure 13 and Figure 14, we validate that HENS has comparable reliability and spread-error ratio as the 58-member ensemble. With a better owCRPS, HENS characterizes the conditional tail distribution better, but it does not come at a cost of overpredicting extremes and obtaining worse performance on the other scores.
@@
Using all forecasts initialized during summer 2023 (Figure 8c), the HENS owCRPS is approximately 20% better than the 58-member ensemble at lead times of 4, 7, and 10 days. This indicates that HENS has a more accurate conditional distribution than that from a smaller 58-member ensemble. The key value-add of HENS is representing these conditional distributions with more fine-grained probabilities. Based on Figure 8b, HENS does not improve the ability to sample the bulk distribution. 58 members are adequate to represent the bulk distribution, and so the 58-member ensemble and HENS have similar CRPS scores. Indeed, ensembles of approximately 50 members have been responsible for the tremendous skill of existing weather forecasts and have created a quiet revolution in numerical weather prediction (Bauer et al., 2015). However, for characterizing the ensemble distribution conditioned on being above a threshold, HENS offers a significant improvement over traditional ensemble sizes.

### v3

@@ 3.1 Sampling the Forecast Distribution with Huge Ensembles
For all statistics, smaller ensemble sizes result in large uncertainties and, in some cases, large biases relative to the full ensemble. Estimates of the HENS mean and standard deviation are unbiased (i.e., the empirical bias is near zero) with as few as 10 members and 200 members, respectively. However, these estimates are associated with large sampling uncertainty, well exceeding six times the uncertainty of the full ensemble. Larger ensembles are necessary for unbiased estimation of the 10 and 90 percentiles, on the order of . For ensembles smaller than 1000, the sign of the bias is notable: estimates of the 10 and 90 percentile are too large and too small, respectively, highlighting the under-sampling of even these moderate percentiles for smaller ensembles. Even with 1000 members, there is still sampling uncertainty associated with moderate percentiles, nearly three times that of the full ensemble, and larger ensembles are necessary to reduce this uncertainty. For the most extreme percentiles, representing 1000-day events, nearly the full ensemble is needed to obtain empirical estimates that are unbiased. In Section G.4.2, we present extreme value methods to calculate the 0.1 and 99.9 percentiles. Compared to directly calculating the percentiles, extreme value theory leads to better estimation of the extremes (Figure 4): estimates are unbiased for ensembles as small as 3000, which is a significant improvement relative to empirical estimates which require . However, as with the 10 and 90 percentiles, for smaller ensembles both the empirical and extreme value theory estimates are not extreme enough, again illustrating that smaller ensembles do not properly sample the extreme tails of the distribution.
@@
Across all statistics considered here, larger ensembles lead to significantly more confident estimation of each property of the full 7,424-member HENS ensemble. These results are robust across different lead times (Figure 18 and for a different variable (see the sampling behavior of 10m wind speed in Figure 19). This is a key value-add of HENS: it enables confident characterization of both the mean and extreme statistics of the forecast distribution, and it quantifies the uncertainty associated with smaller ensemble sizes. Using this information, the users can select a desired ensemble size based on their use case and an acceptable level of uncertainty.
For these analyses (and all future analyses in this manuscript), we note that we do not assume the ensemble distribution is Gaussian. Gaussianity was an emergent property of the global land means, so for these spatial averages, Gaussian theory served as a good estimate of the analytic uncertainty of the ensemble statistics and the information gain. However, at each grid cell, there are significant deviations from Gaussianity (Figure 3). For both global and local forecasts, HENS can robustly sample farther into the tail of the forecast distribution, compared to a 50-member ensemble. In the next sections, we empirically assess the utility of huge ensembles for weather forecasts and for calculating extreme statistics, and we do not make assumptions about the shape of the distributions.
@@
We present a demonstration of the ensemble forecasts at a specific location during a heatwave in the USA Midwest. Kansas City, Missouri, USA had a significant heat-humidity event on August 23, 2023 at 18:00 UTC. According to ERA5, the 2m temperature and dewpoint reached 307 K and 298 K, respectively. To consider the combined effect of both temperature and humidity at the surface, we calculate the heat index introduced by Lu and Romps (2022), which updates the heat index presented in Steadman (1979) to account for particularly hot and humid events. At this time, the heat index in Kansas City was 316 K (43 C).
@@
Figure 5a shows the HENS forecasts of the heat index, as a function of initial time. Despite the significantly larger size of HENS, its ensemble spread becomes narrower with lead time. This indicates that HENS is not overpredicting extreme values at all lead times; its forecasts still have coherent spread as a function of lead time. At a 10-day lead time, both IFS and HENS predicted a warmer than average temperature. However, the verification air temperature and dewpoint temperature lie at the tails of both ensembles’ forecast distributions (Figure 5b). At a ten-day lead time, none of the IFS members successfully capture the magnitude of both 2m air temperature and 2m dewpoint temperature. However, HENS does include members that capture the simultaneous intensity of both these values (Figure 5b). With the large sample size from HENS, it would be possible to study the precursors, drivers, and statistics of the observed extreme. Figure 5 uses the same method to visualize large ensembles as Li et al. (2024).
@@
HENS also enables exploration of counterfactual realities. Some HENS members projected that the heatwave could have been warmer and drier (Figure 5b). Other HENS members indicate that the heatwave may have not occurred altogether, as they predict 2m temperatures near or below the climatological mean. These different counterfactuals can result in different climate impacts: hot and humid extremes can be particularly challenging for human health, while hot, dry extremes can have adverse impacts on crop yields and create conditions conducive to wildfire spread. IFS does not contain any sampling of the counterfactual hot, dry version of the event beyond 311 K air temperature and 291 K dewpoint. We hypothesize that HENS can be used to characterize the dynamical drivers and physical processes associated with each outcome. In particular, HENS could enable a clustering analysis of the drivers and large-scale meteorological patterns that would result in the observed hot, humid heatwave; a hot, dry heatwave; or no heatwave at all. If HENS is reliable and well-calibrated for a given type of extreme, it can also be used to study the conditional probability (given the initial conditions) of different outcomes on medium-range weather time scales.
@@ 3.2 Sampling the Observed Distribution with Huge Ensembles
At each grid cell and time in summer 2023, we convert the ERA5 value into its Z score. Figure 6 shows that HENS provides large sample sizes for the ERA5 events that occurred in summer 2023. In the majority of cases, HENS includes multiple ensemble members that simulate an event that is at least as extreme as the verification value in ERA5. For even the rarest events that are 4 standard deviations away from their climatological mean, HENS includes at least samples of events with at least that magnitude. For events that are 2 and 3 standard deviations away, there are hundreds of ensemble members in HENS that meet or exceed the ERA5 value. We note that there are instances where HENS misses an event entirely: these instances correspond to the whiskers of each event having 0 samples. We quantify this occurrence for heat extremes in Section 4.
@@
The second requirement for using HENS to study rare observed events is that at least some ensemble members should accurately simulate the true event. With these accurate members, it would be possible to study the event’s likelihood and the physical drivers, in comparison to counterfactual outcomes. To assess the suitability of HENS for these types of analyses, we calculate the RMSE of the best ensemble member. Buizza and Palmer (1998) originally introduced this metric for ensembles with up to 32 members. At each grid cell, they choose the ensemble member with the smallest RMSE across all members. Using multiple initialized forecasts and lead times, they quantify the RMSE of this best member.
@@
This metric is actively used in the study of extreme weather events to identify possible drivers. In an operational ensemble weather forecast, Mo et al. (2022) identify the members that had the most accurate forecasts of the 2021 Pacific Northwest Heatwave. These best members correctly forecast the extent and inland location of a warm-season atmospheric river, which served as a source of latent heat for the heatwave. Leach et al. (2024) also examine the ensemble member that is nearest to the observed temperatures. They show that ensemble members that predicted warmer temperatures were associated with low cloud cover and a high geopotential height anomaly. For other types of extreme weather, Millin and Furtado (2022) identify the ensemble members that most accurately simulated a cold air outbreak, and they show that these members correctly forecast two wave breaks, which were dynamical drivers of the event. With HENS, these types of analyses can be conducted at scale.
HENS reduces the RMSE of the best ensemble member; with larger ensembles, this metric systematically decreases by up to 50% (Figure 7). Compared to smaller ensembles, HENS more thoroughly covers the space of possible future outcomes. In this metric, the HENS improvement is greater for more extreme observed events, indicating the benefits of using HENS to study LLHIs. These results are robust across lead time (Figure 20) and for a different variable (Figure 21). This opens the door for future analysis to determine why these members correctly forecast the event and to validate whether they forecast the right value for the right reasons.
We note that the minimum ensemble RMSE metric is not useful in making weather forecasts; while huge ensembles help ensure that at least one member will reasonably match the observations, there is no way of knowing ahead of time which member that will be.
@@ 4 Validating Huge Ensemble Weather Forecasts
4.1 Metrics based on the entire distribution and metrics based on the conditional distribution
Next, we discuss opportunities to improve extreme weather forecasts with larger ensembles. We compare the 58-member ensemble in Part I to the 7,424-member HENS. Even though HENS includes more detailed sampling of the forecast distribution, its overall CRPS and threshold-weighted CRPS (twCRPS) scores do not change significantly. We provide a theoretical basis and a case study to understand the invariance of these two scores with ensemble size. These scores are calculated using the entire forecast distribution, but HENS is beneficial for sampling the conditional distribution, given that the forecast is greater than a climatological percentile. We quantify this benefit through the outcome-weighted CRPS (owCRPS).
On the CRPS and twCRPS scores, HENS performs slightly better than the 58-member ensemble, but the improvement is less than approximately five percent (Figure 8a and b). We note that we are not using versions of CRPS that are debiased with respect to ensemble size (Zamo and Naveau, 2017). The twCRPS is calculated analogously to CRPS except all values of the ensemble forecast below a pre-specified threshold are converted to the threshold itself (Allen et al., 2023). As a function of ensemble size, these scores appear to have largely converged with 58 members.
@@ 4.1 Metrics based on the entire distribution and metrics based on the conditional distribution
CRPS and twCRPS compare the CDF of the ensemble forecast to the verification value (see Equations 1 and 3 of Part I). The CDFs are constructed from all members of the ensemble. The Dvoretzky–Kiefer–Wolfowitz (DKW) inequality (Massart, 1990) quantifies the maximum difference between a true population CDF and an empirical CDF constructed from samples. Here, the HENS forecast is the population CDF, and the 58-member ensemble is the sample empirical CDF. The difference between these two CDFs scales with a factor of (Appendix Section H). With this scaling factor, a 58-member CDF closely approximates a 7,424-member CDF, so we expect the corresponding CRPS and twCRPS to be very similar. Based on this argument, we also expect the Extreme Forecast Index, which is also calculated from a CDF based on the entire ensemble (see equation 2 of Part I), to be very similar in HENS and the 58-member ensemble.
@@
However, HENS has an advantage in resolving the conditional distribution, given that the ensemble is above a climatological threshold. This conditional distribution is created by truncating the ensemble forecast distribution at the climatological 95th percentile. When the CRPS is calculated using the conditional distribution, it is referred to as the outcome-weighted CRPS (owCRPS). Mathematically, owCRPS is
@@
where is a random variable from the ensemble CDF. We use the weighing function , based on a threshold , which is the 99th percentile of 24-year ERA5 climatology as the threshold.
@@
In Figure 9a, we show a case study of the full ensembles’ CDFs at Shreveport, Louisiana, USA. Due to its smaller size, there is more noise in the 58-member CDF compared to the HENS CDF. In Figures 9b, c, and d, we compare the distributions that are used to calculate CRPS, twCRPS, and owCRPS. These three scores use the raw ensemble distribution, the transformed ensemble distribution, and the conditional distribution, respectively. twCRPS results in a point mass being placed at the threshold (Figure 9c), since all ensemble members below the threshold are converted to having a value at the threshold itself. owCRPS is created from the conditional distribution. Figures 9b, c, and d mirror the schematic in Figure 1 of Allen et al. (2023), which demonstrates calculating CRPS, twCRPS, and owCRPS.
@@
In this case study, HENS shows small improvements in CRPS and twCRPS, which use the entire ensemble distribution, but it shows a large improvement in the owCRPS, which uses the conditional distribution. In Figure 9d, the 58-member ensemble had 10 members above the threshold, and HENS had 1340 members above the threshold. Increasing the ensemble size from 10 to 1340 members yields a greater sampling improvement than increasing the ensemble size from 58 to 7424 members. In this case study, 10 members are not enough to adequately characterize the tail of the forecast distribution (Figure 9d). Because of the significant sampling uncertainty, the small ensemble has a different conditional distribution than HENS’s conditional distribution. HENS reduces the owCRPS from 0.63 K to 0.5 K (a relative reduction of about 20%). Unlike CRPS and twCRPS, the number of ensemble members used in the owCRPS varies for each forecast at each grid cell: it depends on how many members are above the threshold.
@@
The owCRPS is calculated only when the extreme event actually occurs, so it is not a statistically proper scoring rule. Proper scoring rules are minimized when the forecast distribution matches the distribution from which the observation is drawn (Gneiting and Raftery, 2007). They cannot be hedged by overpredicting extremes (see Part 1 for a discussion on the relationship between CRPS and proper scoring rules for extreme forecasts). Since the owCRPS is only calculated when extremes actually occur, a forecast that overpredicts extremes could falsely appear reliable (Lerch et al., 2017). This is the essence of the forecaster’s dilemma (see Part I). However, HENS does not appear to be overpredicting extremes and hedging its scores because it has a comparable CRPS, twCRPS, reliability, and spread-error ratio as the 58-member ensemble. If HENS were overpredicting extreme weather, then these scores would degrade. In Figure 13, Figure 14, Figure 15, we also show that HENS has comparable reliability, spread, and error as the 58-member ensemble.
@@
Using all forecasts initialized during summer 2023 (Figure 8c), the HENS owCRPS is approximately 20% better than the 58-member ensemble at lead times of 4, 7, and 10 days. Based on Figure 8b, HENS moderately improves the ability to sample the bulk distribution (e.g. through the reduced noise in Figure 9), and this could result in the small improvements to CRPS and twCRPS. In many cases, a 58-member ensemble is adequate to represent the bulk distribution. Indeed, ensembles of approximately 50 members have been responsible for the tremendous skill of existing ensemble weather forecasts. To characterize the ensemble distribution conditioned on exceeding a threshold, only a subset of members can contribute to the conditional CDF. In this scenario, HENS provides a significant advantage over traditional ensemble sizes.